If you are going 20 mph., how many inches per second are you traveling?




If you are going 72 mph., how mnay feet per minute are you traveling?

Answer by jhadipak(3)

Q.1: If you are going 20 mph., how many inches per second are you traveling?
Solution: we have been given the speed as 20 mph.
we need to find the speed in inch per seconds.
we know that 1 mph = 17.6 inches per seconds
so, 20 mph = 20(17.6) inches per seconds
= 352 inches per seconds
Or, I am going to explain how 1 mph = 17.6 inches per seconds
we know that 1 mile = 63360 inches
and per hour = 60 minutes
per minute = 60 seconds
so, in one hour = 60*60 seconds = 3600 seconds
1 mph = [1*(63360)]/3600 = 17.6 inch per seconds

Q.2:If you are going 72 mph., how many feet per minute are you traveling?
Solution: we know that 1 mph = 88 feet per minutes
so for 72 mph = 72 * 88 feet per minutes
= 6336 feet per minutes.
Hope this will help you!!!